# Machine Unlearning Doesn’t Do What You Think: Lessons for Generative AI Policy, Research, and Practice

### v1

@@ Abstract
We articulate fundamental mismatches between technical methods for machine unlearning in Generative AI, and documented aspirations for broader impact that these methods could have for law and policy. These aspirations are both numerous and varied, motivated by issues that pertain to privacy, copyright, safety, and more. For example, unlearning is often invoked as a solution for removing the effects of targeted information from a generative-AI model’s parameters, e.g., a particular individual’s personal data or in-copyright expression of Spiderman that was included in the model’s training data. Unlearning is also proposed as a way to prevent a model from generating targeted types of information in its outputs, e.g., generations that closely resemble a particular individual’s data or reflect the concept of “Spiderman.” Both of these goals—the targeted removal of information from a model and the targeted suppression of information from a model’s outputs—present various technical and substantive challenges. We provide a framework for thinking rigorously about these challenges, which enables us to be clear about why unlearning is not a general-purpose solution for circumscribing generative-AI model behavior in service of broader positive impact. We aim for conceptual clarity and to encourage more thoughtful communication among machine learning (ML), law, and policy experts who seek to develop and apply technical methods for compliance with policy objectives.
Authors’ expertise and intended audience.  Our contributions require expertise in ML, law, and policy. We intend for our audience to be members of all of those communities. We organized a team of experts in each discipline. The resulting paper reflects the efforts of a large-scale collaboration across academic institutions, civil society, and industry labs. We intend for this paper to be a standalone document: one with the necessary (and sometimes elementary) background to make our contributions legible to our diverse intended audience, and at the appropriate level of abstraction to encourage effective cross-disciplinary communication about machine unlearning.
@@ 1 Introduction
“Machine unlearning” has recently captured public attention as a potential general-purpose approach for purging unwanted information from machine-learning (ML) models. It raises problems of technical interest, but perhaps more significantly, machine unlearning also finds broader appeal outside of technical circles for its perceived ability to support law and policy aims (Section 2). Since around 2016, technical experts and policymakers have invoked unlearning as a way to operationalize compliance with an individual’s “right to be forgotten,” with respect to removing personal data from deployed models, as granted in the E.U.’s General Data Protection Regulation (GDPR) [102]. [^1] This paper focuses on generative-AI models, but unlearning methods originate from classification. Now, with the emergence of Generative AI, machine unlearning’s presumptive mandate has expanded significantly. More and more, research papers, policy briefs, and media reports suggest machine unlearning as an approach for meeting a broad range of objectives for both open and closed models and systems, [^2] Our observations address cross-cutting issues applicable to both open and closed technology. spanning privacy, copyright, safety, and more [e.g., 134, 71, 63, 75, 51, 118, 96, 140].
Unfortunately, the fit between unlearning and policy is not so straightforward in practice. Machine unlearning is a set of technical methods and here, as always, there are critical gaps—gaps that are too often overlooked—between what technical methods do and what policy aims to achieve [31]. Our goal is to provide conceptual clarity that elicits these gaps, and to encourage more thoughtful communication among ML, law, and policy experts who seek to develop and apply technical methods for compliance with law and policy objectives (Sections 6 & 7). In summary:
@@
Deleting information from an ML model is not well-defined.  First, information cannot be deleted from an ML model in the same way that it can from a database. During training, data are transformed into patterns that get encoded in the model’s parameters—patterns that are not directly or easily interpretable (Section 2). There is no way to cleanly identify, target, and delete specific, contained pieces of information from these parameters. Instead, it is possible, even if computationally expensive, to train a new model on a dataset that does not contain problematic data (Section 4)—for example, a specific scientific paper on designing novel flu viruses or a specific in-copyright image of Spiderman. This is typically what it means to “remove” data from a generative-AI model in machine unlearning, which deviates from intuitive understandings of the term. Removal applies to discrete pieces of data in the training dataset before training occurs; it cannot target the latent patterns that a trained model has learned across different data examples (Section 3). For example, there is no clear way to remove the more general concepts of “how to synthesize a toxic molecule” or “Spiderman” from a model; there is no single obvious or appropriate way to go about translating such open-ended aims to concrete tasks that can be implemented by an algorithm (Section 5).
@@
Removing information from a model does not provide guarantees about model outputs.  Second, removing information from a model’s parameters does not guarantee that this model could never produce related information at generation time. Even if one removed all in-copyright images of Spiderman from a model’s training data, this does not mean it would be impossible for the model to generate outputs that resemble Spiderman when put to use. Generative-AI models are impressive in part because they are able to generate novel outputs that transcend the information that is exactly contained in their training data. It is therefore a mistake to think that making a limited set of targeted changes to a model’s parameters is sufficient to make promises about what types of outputs that model could or could not possibly produce (Section 5). This point is further complicated by the fact that users can introduce information at generation time through prompts. In the context of machine unlearning, user prompts can even reintroduce information whose effects were previously removed from the model’s parameters. Combining such a prompt with a model’s reasoning abilities, it may be possible to produce outputs that are effectively the same as those that would have been produced if an unlearning method had never been used in the first place (Section 5).
In general, removal on its own is often neither necessary nor sufficient to constrain model outputs in a controlled manner. [^3] One important exception where removal may be necessary (but still not sufficient) concerns illegal content that may have been included in the model’s training data, such as child sexual abuse material (CSAM) [101]. Instead, suppressing certain model outputs from being surfaced to users may be a more appropriate area of focus for technical methods [e.g., 39]. While it is now common to include output suppression under the umbrella of “machine unlearning,” arguably, these methods have nothing to do with “unlearning” some information from a model; they serve as guardrails on model and system outputs that bear more resemblance to alignment techniques (Section 4).
@@
Even seemingly innocuous model outputs can be put to undesirable uses.  Lastly, even if technical research were to shift the focus to methods for suppressing undesirable model outputs, this would not immediately provide solutions for all law and policy ends. Generative-AI systems are dual-use, where the appropriateness of downstream use wholly depends on context. There remain familiar (but fundamentally irresolvable) tensions that are inherent to highly generative, dual-use technologies [151], like the PC and the Internet (Section 7). Just as a PC could be used as a tool to perpetrate fraud or to write the next great Broadway musical, a generative-AI system can similarly be put to malicious or beneficial uses. Further, on their own, generated outputs may be innocuous or have significant legitimate uses; yet these same outputs could be pressed into service for adversarial or malicious downstream uses. To greater or lesser extents, different unlearning methods can remove the influence of specific training data from a model’s parameters or suppress undesirable model outputs (Section 4). But the type of control these functions provide is localized to the model or system. Additional controls on downstream use would require anticipating how a person or other agent might behave with outputs in a potentially infinite number of contexts—none of which is reasonably under the purview of machine unlearning (Section 5).
In this paper, we explore these observations in detail, and examine their various implications for privacy, copyright, and safety—three areas for which machine unlearning has been suggested as a viable approach for operationalizing compliance with law or policy requirements (Section 6). Rather than following the well-trod path of surveying [109, 143, 83, 117, 114, 82] or evaluating [81, 83, 89] existing unlearning methods, we take a step back and think conceptually about what, in principle, machine unlearning could reasonably accomplish.
We articulate fundamental mismatches between machine unlearning—as a technical problem of study in ML research—and aspirations for the broader impact that methods emerging from this research could have for law and policy. In contrast to common opinions in policy research [e.g., 9, 63], we show that machine unlearning—both the entire class of methods and specific techniques—should not be misunderstood as a general-purpose solution for circumscribing model behavior in service of broader positive impact. Unlearning methods are imperfect and may serve as only one approach of many that could, in some cases, contribute to addressing aspects of issues that are of interest to policymakers. However, given the fundamental mismatches we identify, it is also hard to imagine that, even with time, technical solutions for unlearning will ever wholly achieve desired law and policy objectives. In light of these limitations, we provide recommendations on how ML experts should focus their research and how policymakers can adjust their expectations and norms concerning reasonable best efforts when using an unlearning method in practice (Section 7).
We organize the remainder of the paper as follows:
@@
We begin with some necessary background on machine unlearning—both its technical motivations and evolving motivations for Generative AI from law and policy.
@@
We identify different targets—observed information, latent information, and higher-order concepts—that model developers or custodians may want to address with unlearning.
@@
Defining these targets (Section 3) helps us make clear which types of information a specific unlearning method may apply to and which it does not. Some methods can remove targeted pieces of observed information before training occurs. For Generative AI, most methods aim to suppress model outputs that contain undesirable content.
@@
Together, our discussion illuminates four important mismatches between unlearning motivations (Section 2), targets (Section 3), and methods (Section 4). These mismatches lay the groundwork for understanding how there are substantive aims that cannot, from first principles, be addressed with unlearning methods alone.
@@
We examine how these mismatches (Section 5) manifest differently and exhibit various implications for issues related to privacy, copyright, and safety contexts.
Section 7.
We suggest takeaways and possible future directions for ML research and AI policy.
@@ 2 Background and Motivations for Machine Unlearning
The natural starting place for our discussion is to address first what machine learning attempts to accomplish. This will let us provide an intuition, grounded in an interpretation of the E.U.’s General Data Protection Regulation (GDPR) that is prevalent in ML research, for why one might want to do unlearning to revert or change the results of this process (Section 2.1). From this intuition, we provide a loose definition for machine unlearning that originates from traditional AI settings (Section 2.2). We then discuss evolving motivations for machine unlearning in response to the ascendance of Generative AI (Section 2.3). These new motivations have encouraged an expanded definition for machine unlearning (Section 2.4), which we will rely on throughout the paper.
2.1 ML research and its interpretation of the GDPR
In brief, machine learning is an area of computer science and engineering that uses techniques from probability and statistics to develop algorithms that produce models that encode patterns learned from data. Model architectures range from simple linear models or decision trees to complicated, large-scale neural networks. In a bit more detail, we rely on the GenLaw glossary [31]: [^4] We reprint with permission excerpts from the GenLaw Glossary. This glossary provides definitions in machine learning, law, and policy at the same level of abstraction that is intended for our audience.
  • Machine-learning neural-network models all contain parameters. …During an algorithmic process called training, these parameters are repeatedly updated based on the training data within the training dataset that the model has seen. Each update is designed to increase the chance that when a model is provided some input, it outputs a value close to the target value we would like it to output. By presenting the model with all of the examples in a dataset and updating the parameters after each presentation, the model can become quite good at doing the task we want it to do.
Each training-data example in the training dataset “is a self-contained piece of data, such as an image, a piece of text (e.g., content of a web page), a sound snippet, a video, or some combination of these” [31]. These examples can also include personal information—home addresses, sensitive demographic attributes, health information, personal photos, and more.
@@ 2.1 ML research and its interpretation of the GDPR
In some jurisdictions, individuals have rights associated with the control of their personal data. Notably, since its adoption in 2018, Article 17 of the E.U.’s GDPR provides the “Right to erasure” (more commonly called the “right to be forgotten”) [102, 45], which gives individuals rights (with exceptions) to demand that companies delete their personal data. [^5] There are several exceptions to this right, for instance, when keeping the data is in the public interest (e.g., the data are used to comply with a legal ruling) or for certain research purposes, etc. [102, Article 17(3)]. ML researchers often interpret Article 17 to apply to both to the individual’s data examples that have been used as training data and to the resulting trained models themselves [e.g., 73, 84, 24, 90, 106, 74, 10] (Appendix A). [^6] While this interpretation is common in unlearning papers, it is still very much up for debate. Exactly how Article 17 may or may not apply to ML models is under active discussion. (See Section 6.1.) This presents a problem because, in almost all cases, the model would not just be trained on a specific, right-exercising individual’s data. It would also be trained on data associated with thousands of others, if not many more. Wholesale erasure of a trained model, in response to one individual’s deletion request, would therefore likely be an extreme, over-broad interpretation of Article 17.777Also consider that many thousands of people might make such a request, each one requiring retraining a new model from scratch (Section 4.1). Retraining runs could perhaps be periodically batched—removing multiple individuals’ personal data together in the same run to reduce the number of times a model is retrained. Even still, retraining could happen a large number of times for the same model.
@@
This problem raises a natural question for ML research: rather than deleting a trained model altogether, is it possible to develop algorithms that can achieve more targeted removal of training data from the model? In the specific context of ML research’s common interpretation of the GDPR: is it possible to remove the influence of the right-exercising individual’s data from the model, without imposing on the model controller the undue burden of the cost of retraining a new model from scratch without that individual’s data examples (Section 4.1)? Machine unlearning is the area of ML research that attempts to address this question.
2.2 A loose definition for machine unlearning
What is referred to as “machine unlearning” in the technical literature actually corresponds to a wide variety of different methods and techniques, which are loosely grouped together. For this reason, we will rely on a loose, intuitive (rather than rigorous) definition of machine unlearning that we construct in relation to this common underlying technical motivation. In relation to early research, machine unlearning is a subarea of machine learning that develops methods for the targeted removal of the effect of training data from the trained model. We will soon refine this definition (Section 2.4), in response to changing motivations in the field for when to use unlearning (Section 2.3) and how to implement it (Section 4).

### v2

# Machine Unlearning Doesn’t Do What You Think: Lessons for Generative AI Policy and Research

@@ Abstract
“Machine unlearning” is a popular proposed solution for mitigating the existence of content in an AI model that is problematic for legal or moral reasons, including privacy, copyright, safety, and more. For example, unlearning is often invoked as a solution for removing the effects of specific information from a generative-AI model’s parameters, e.g., a particular individual’s personal data or the inclusion of copyrighted content in the model’s training data. Unlearning is also proposed as a way to prevent a model from generating targeted types of information in its outputs, e.g., generations that closely resemble a particular individual’s data or reflect the concept of “Spiderman.” Both of these goals—the targeted removal of information from a model and the targeted suppression of information from a model’s outputs—present various technical and substantive challenges. We provide a framework for ML researchers and policymakers to think rigorously about these challenges, identifying several mismatches between the goals of unlearning and feasible implementations. These mismatches explain why unlearning is not a general-purpose solution for circumscribing generative-AI model behavior in service of broader positive impact.
@@ 1 Introduction
Since around 2016, technical experts and policymakers have invoked machine unlearning as a way to operationalize compliance with an individual’s “right to be forgotten” in the E.U.’s General Data Protection Regulation (GDPR) [111], with respect to removing personal data from deployed models. More recently, with the emergence of generative AI, machine unlearning has captured public attention as a potential general-purpose approach for purging unwanted information from generative-AI models and systems. More and more, research papers, policy briefs, and media reports suggest that unlearning can help meet a broad range of objectives for both open and closed models and systems, [^1] Our observations address cross-cutting issues applicable to both open and closed generative-AI technology—and technologies that lie on the spectrum between these two poles. spanning privacy, copyright, safety, and other issues [e.g., 140, 76, 84, 54, 126, 105, 146, 9, 69].
For generative AI, the goal of unlearning has shifted beyond the traditional ML focus on removing the influence of specific training data from model parameters. This kind of targeted removal is often infeasible in this setting, and even when possible, it may not prevent problematic content from appearing in a model’s outputs. [^2] One important area where removal may be necessary (but still not sufficient) concerns illegal or otherwise prohibited content that may have been included in the model’s training data, such as child sexual abuse material (CSAM) [110]. See Section 5. As a result, unlearning now also includes targeted suppression at the output level, reflecting an attempt to manage what generative-AI models produce, not just what is encoded in their parameters. But these are fundamentally different goals—both in the technical methods they require (Section 3) and in how their outcomes may be treated under specific legal provisions (Section 5). As a result, mapping unlearning methods onto policy objectives is not straightforward; important mismatches and slippages arise, for three overarching reasons:
@@
Deleting information from an ML model is not like deleting data from a database.  There is no way to cleanly identify, target, and delete specific, contained pieces of information from an ML model’s parameters. Instead, it is possible (though expensive) to train a new model on a dataset that does not contain problematic data (Section 3)—e.g., a specific scientific paper on designing novel viruses or a specific in-copyright image of Spiderman. This is typically what it means to “remove” data from a model in unlearning, which deviates from intuitive understandings of the term. Removal applies to discrete pieces of data in the training dataset before training occurs; it cannot target latent patterns a trained model has learned across different data—e.g., more general concepts of “how to synthesize a toxic molecule” or “Spiderman.” There is also no obvious or appropriate way to go about translating such open-ended aims to concrete tasks that can be implemented by an algorithm (Section 4).
@@
Removing information from model parameters does not give guarantees about model outputs.  It is possible to retrain a model so that specific data are removed from the training dataset, and thus those removed data do not influence the model’s parameters. But doing so will not necessarily prevent the model from producing outputs that resemble removed data. For instance, even if one removed all in-copyright images of Spiderman from a model’s training data, this does not mean it would be impossible for that model to produce outputs that resemble Spiderman at generation time. Generative-AI models are impressive in part because they generalize beyond the information that is exactly contained in their training data. It is therefore a mistake to think that making a limited set of targeted changes to a model’s parameters is sufficient to make promises or guarantees about what types of outputs that model could or could not possibly generate (Section 4).
@@
Suppression methods are necessarily imperfect.  While it is technically feasible to suppress certain types of outputs, suppression techniques (e.g., output filters) are unlikely to produce perfect legal compliance. This is in part because generative-AI systems exhibit familiar but fundamentally irresolvable tensions that are inherent to highly generative technologies [158], like the PC and the Internet (Section 6). Just as a PC could be used as a tool to perpetrate fraud or to write the next great Broadway musical, a generative-AI system can similarly be put to malicious or beneficial uses. Suppression methods cannot control downstream use; this would require anticipating how a person or other agent might behave with outputs in a potentially infinite number of contexts—none of which is reasonably under the purview of a technical method alone (Section 4).
We explore each of these points in detail. Altogether, we articulate fundamental, conceptual mismatches between technical methods in machine unlearning (Section 4) and aspirations for the broader impact that these methods could have for law and policy (Section 57). [^3] Our contributions require expertise in ML, law, and policy. We intend for our audience to be members of all of those communities. We organized a team of experts in each discipline. The resulting paper reflects the efforts of a large-scale collaboration across academic institutions, civil society, and industry labs. We intend for this paper to be a standalone document: one with the necessary (and sometimes elementary) background information to make our contributions legible to our diverse intended audience, and at the appropriate level of abstraction to encourage effective cross-disciplinary communication about machine unlearning. While machine unlearning research will continue to progress, we are doubtful that unlearning methods will deliver full compliance in the future. Nevertheless, unlearning techniques may offer limited benefits that support certain outcomes for law and policy. In summary:
@@
We discuss evolving motivations for machine unlearning in generative AI—to extend beyond the targeted removal of the influence of specific training data on a model’s parameters to also include the targeted suppression of specific content from model outputs.
@@
We briefly discuss common technical approaches for both removal and suppression.
@@
We show how our discussion illuminates five conceptual mismatches between unlearning motivations and concrete technical methods—mismatches that make clear there are substantive aims that cannot, from first principles, be addressed with unlearning methods alone.
@@
We examine how these mismatches manifest differently and exhibit various implications for U.S. copyright law, privacy, and safety.
@@
In light of these limitations, we provide recommendations on how ML experts should focus their research and how policymakers can adjust their expectations and norms concerning reasonable best efforts when relying on an unlearning method in practice.
@@ 2 Background and motivations for machine unlearning
Early motivations for supervised machine unlearning in the GDPR.  In some jurisdictions, individuals have rights associated with the control of their personal data. Notably, since its adoption in 2018, Article 17 of the E.U.’s GDPR provides the “Right to erasure” (more commonly called the “right to be forgotten”) [111, 47], which gives individuals broad rights (with exceptions) to demand that companies delete their personal data. [^4] There are several exceptions to this right, for instance, when keeping the data is in the public interest (e.g., the data are used to comply with a legal ruling) or for certain research purposes, etc. [111, Article 17(3)]. ML researchers often interpret Article 17 to apply to both to the individual’s data examples that have been used as training data and to the resulting trained models themselves [e.g., 82, 92, 24, 98, 114, 83, 10] (Appendix A). [^5] While this interpretation is common in unlearning papers, it is still very much up for debate. Exactly how Article 17 may or may not apply to ML models is under active discussion. (See Section 5.2.) This presents a problem because, in almost all cases, the model would not just be trained on a specific, right-exercising individual’s data. It would also be trained on data associated with thousands of others, if not many more. Wholesale erasure of a trained model, in response to one individual’s deletion request, would therefore likely be an extreme, over-broad interpretation of Article 17.666Also consider that many thousands of people might make such a request, each one requiring retraining a new model from scratch (Section 3.1). Retraining runs could perhaps be periodically batched—removing multiple individuals’ personal data together in the same run to reduce the number of times a model is retrained. Still, retraining could happen a large number of times for the same model.
@@
This problem raises a natural question for ML research: rather than deleting a trained model altogether, is it possible to develop algorithms that can achieve more targeted removal of training data from the model? In the specific context of ML research’s common interpretation of the GDPR: is it possible to remove the influence of the right-exercising individual’s data from the model, without imposing on the model controller the undue burden of the cost of retraining a new model from scratch without that individual’s data examples (Section 3.1)?